OpenStudy (anonymous):

(5√2-15√5)(4√2+12√5) I got (√27+√230)(√18+√149) √486+√4023+√4140+√34,270 Is this the final answer or can it be simplified...

OpenStudy (asnaseer):

\[\begin{align} (5\sqrt{2}-15\sqrt{5})(4\sqrt{2}+12\sqrt{5})&=(5\sqrt{2})(4\sqrt{2})+(5\sqrt{2})(12\sqrt{5})\\ &-(15\sqrt{5})(4\sqrt{2})-(15\sqrt{5})(12\sqrt{5})\\ &=(20*\sqrt{2}*\sqrt{2})+(60*\sqrt{2}*\sqrt{5})\\ &-(60*\sqrt{5}*\sqrt{2})-(180*\sqrt{5}*\sqrt{5})\\ &=(20*2)+(60*\sqrt{10})-(60*\sqrt{10})-(180*5)\\ &=(40)-(900)\\ &=-860 \end{align}\]